OpenStudy (anonymous):

Can someone please help me : You wish to invest $90,000 in two accounts. You want a very aggressive investment strategy, so you intend to invest in an account with a rate of 11.8% and another with a rate of 22.3%. With this strategy, at the end of one year, you want to earn $16,668.00. How much should you invest in each account?

OpenStudy (anonymous):

System of linear equations

OpenStudy (anonymous):

Let money in one account be 'x' and money in other account be 'y' This gives you your first equation, x + y = 90000. This goes to: y = 90000 - x Let the interest rate of account where you invest 'x' be 11.8%. The return on this account will by 1.118x Let the interest rate of account where you invest 'y' be 22.3%. The return on this account will be 1.223y This gives you your second equation, 1.118x + 1.223y = 106668 Put y = 90000 -x in equation above. That will give you x= 32400, and hence y=57600

OpenStudy (anonymous):

thank you!